# Taming Data and Transformers for Audio Generation

### v2

@@ Abstract
Generating ambient sounds is a challenging task due to data scarcity and often insufficient caption quality, making it difficult to employ large-scale generative models for the task. In this work, we tackle this problem by introducing two new models. First, we propose AutoCap, a high-quality and efficient automatic audio captioning model. By using a compact audio representation and leveraging audio metadata, AutoCap substantially enhances caption quality, reaching a CIDEr score of $83.2$, marking a $3.2\%$ improvement from the best available captioning model at four times faster inference speed. Second, we propose GenAu, a scalable transformer-based audio generation architecture that we scale up to 1.25B parameters. Using AutoCap to generate caption clips from existing audio datasets, we demonstrate the benefits of data scaling with synthetic captions as well as model size scaling. When compared to state-of-the-art audio generators trained at similar size and data scale, GenAu obtains significant improvements of $4.7\%$ in FAD score, $22.7\%$ in IS, and $13.5\%$ in CLAP score, indicating significantly improved quality of generated audio compared to previous works. Moreover, we propose an efficient and scalable pipeline for collecting audio datasets, enabling us to compile 57M ambient audio clips, forming AutoReCap-XL, the largest available audio-text dataset, at 90 times the scale of existing ones. Our code, model checkpoints, and dataset are publicly available.
@@ 1 Introduction
Text-conditioned generative models have revolutionized the field of content creation, enabling the generation of high-quality natural images (Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023; Haji-Ali et al., 2024), vivid videos  Ho et al. (2022); Villegas et al. (2022); Wang et al. (2023b); Qiu et al. (2023); Menapace et al. (2024), and intricate 3D shapes (Cheng et al., 2023). The domain of audio synthesis has undergone comparable advancement (Huang et al., 2023b; a; Liu et al., 2023b; Xue et al., 2024; Guan et al., 2024; Saito et al., 2024; Niu et al., 2024; Yang et al., 2023a; Evans et al., 2024b; Liu et al., 2024; Wang et al., 2024b; Guo et al., 2023), with three broad areas of study: speech, music and ambient sounds. The success in these domains rests on two key pillars: (i) the availability of high-quality large-scale datasets with text annotations, and (ii) the development of scalable generative models (Ho et al., 2020; Song et al., 2020).
@@
In the field of audio synthesis, ambient audio generation emerges as a critical domain, which is the main focus of this work. Unlike speech and music, ambient sound generation suffers from a lack of extensive, well-annotated datasets (Kim et al., 2019; Drossos et al., 2020). Attempts to curate ambient audio from online videos predominantly failed, primarily due to the dominance of speech and music content in such videos. For instance, AudioSet (Gemmeke et al., 2017), the largest available audio dataset sourced from online videos, contains $99\%$ speech or music clips. Previous efforts to filter ambient audio from similar datasets involved using expensive classifiers on the video or audio content, making it impractical to compile a large-scale dataset due to the high rejection rate. In this work, we propose a simple, yet scalable filtering approach that leverages existing automatic video transcription to identify segments with ambient sounds. This method is not only more efficient but also more feasible, as it eliminates the need to download the audio or video content. Additionally, by using time-aligned transcripts, we reduce the rejection rate to only $83\%$. Through this approach, we built AutoReCap-XL, a dataset containing 57 million ambient audio clips sourced from existing video datasets, representing a 90-fold increase over the size of previously available datasets.
@@
Another challenge in compiling large-scale text-audio datasets is providing accurate textual descriptions. For visual modalities, such as images and videos (Xue et al., 2022; Miech et al., 2019), researchers often relied on the raw description and metadata to train strong visual-text models including reliable captioners (Chen et al., 2024b). Similarly, speech modality benefits from strong automatic transcription models to provide textual annotations. For ambient sounds, however, the task is substantially more challenging as accompanying raw text tends to describe visual information or convey feelings, rather than detailing the audio content. Moreover, human-captioned audio datasets are limited, containing fewer than $51k$ text-audio pairs in total. This significantly impacts the training of current captioning models, making them more susceptible to overfitting and reducing their ability to generalize effectively. In this work, we address this challenge by introducing AutoCap, an efficient and high-quality audio captioning model that leverages visual information to enhance captioning.
@@
AutoCap refines the commonly used encoder-decoder design based on a pretrained BART (Lewis et al., 2020) model by introducing a Q-Former (Li et al., 2023a) that learns to summarize the encoded audio tokens into four times fewer tokens. By reducing the number of input tokens to the BART model, we speed up inference—an important step towards large-scale audio captioning—and provide better alignment with the original BART token representation due to the Q-Former additional capacity compared to simple projection layers used in previous work (Kim et al., 2024b). Second, we propose to use metadata and captions derived from video content to aid the captioning process and in this way, remedy the data scarcity problem. Critically, we augment the encoder inputs to assume both audio features and a set of descriptive textual metadata including audio title and a caption derived from the visual modality. This dual-input approach not only allows our model to achieve state-of-the-art performance on AudioCaps (Kim et al., 2019), marking a 3.2% improvement in CIDEr score, but it also helps reduce the domain gap with in-the-wild audios.
@@
Moreover, to adapt audio generative models for larger scale training, we introduce GenAu, a scalable transformer-based architecture that achieves significant improvements over state-of-the-art audio generation models. Our approach introduces key architectural modifications over existing audio latent diffusion models (Liu et al., 2023b; Huang et al., 2023b; Ghosal et al., 2023; Huang et al., 2023a). First, we train an efficient 1D-VAE (Huang et al., 2023a) to transform a Mel-Spectrogram representation to a sequence of tokens and search for the optimal latent space for audio generation. Second, we recognize that audio grows fast temporally, yet contains many silent and redundant segments. Therefore, an efficient architecture that can handle such properties is needed. In particular, we employ a transformer architecture in the denoising backbone where we modify the FIT transformer (Chen & Li, 2023) to generate audio in the latent space. Lastly, we extend the proposed FIT architecture to incorporate text conditioning through a dual encoder strategy. This involves an instruction-finetuned language model, FLAN-T5 (Chung et al., 2022), and an audio-centric CLAP encoding (Kim et al., 2024b). This adaptations significantly improves the model’s performance over exiting methods, achieving $22.7\%$ higher Inception Score, $4.7\%$ better FAD, and $13.5\%$ improvement in CLAP score, demonstrating superior audio-text alignment and audio generation quality.
Finally, we explore the scaling behavior of text-to-audio diffusion models in relation to model size and data size. Prior studies in text-to-image generation have established a scaling trend where performance improves with the increases in data volume and model size (Peebles & Xie, 2023; Li et al., 2024). Such exploration, however, has not been sufficiently conducted for the audio modality. Initially, we analyze the impact of augmenting the dataset with synthetic captions on model performance. Our findings reveal a clear trend of improvement in FD and IS as we increase the amount of training data. Furthermore, we observe a consistent trend of enhanced performance across all metrics when scaling up the model size, concluding that the audio modality also benefits significantly from increases in both model size and data scale.
@@
In summary, this work introduces: (i) AutoCap, a state-of-the-art audio captioner tailored towards the annotation of data at a large scale, which leverages audio metadata to improve accuracy and robustness, and a Q-Former to improve inference time and reduce overfitting; (ii) GenAu, a novel audio generator based on a scalable transformer architecture specifically adapted to the audio domain. Our model achieves significantly improved quality when compared to the previous state-of-the-art. (iii) AutoReCap-XL, the largest available audio dataset, comprising 57M audio clips paired with synthetic captions derived from the proposed audio captioner.
@@ 2 Related Work
Automatic Audio Captioning (AAC). The goal of AAC is to produce natural language descriptions for given audio content. Most recent AAC methods (Deshmukh et al., 2023; Wu et al., 2024; Salewski et al., 2023; Sridhar et al., 2023; Kadlčík et al., 2023; Cousin et al., 2023; Labbé et al., 2023; Xu et al., 2023; Zhang et al., 2024) employ encoder-decoder transformer architectures, where an encoder receiving the audio signal produces a representation that is used by the decoder to produce the output caption. WavCaps (Mei et al., 2023a) employs the CNN14 (Kong et al., 2019) and HTSAT (Chen et al., 2022) audio encoders and uses a pretrained BART (Lewis et al., 2020) language decoder. CoNeTTE (Étienne Labbé et al., 2023) proposes an audio encoder based on the ConvNeXt architecture and uses a vanilla transformer decoder (Vaswani et al., 2017) trained from scratch. Recently, EnCLAP (Kim et al., 2024b) proposes the joint use of two audio representations in the form of CLAP (Elizalde et al., 2023) sequence embeddings and a discrete EnCodec (Défossez et al., 2022) audio representation and uses a pretrained BART model as the language backbone. Other work explores augmentation strategies to counter data scarcity (Kim et al., 2022; Étienne Labbé et al., 2023; Ye et al., 2022). Liu et al. (2023d) recently proposed to leverage the visual information using a pre-trained visual encoder to address sound ambiguities, reporting improvements. BART-Tags (Gontier et al., 2021) generates captions conditioned on a sequence of predicted AudioSet tags. Our method uses audio metadata and visual information as additional conditioning signals and leverages a lightweight Q-Former (Li et al., 2023a) model that summarizes the audio feature to improve captioning speed and reduce overfitting.
@@
Text-conditioned audio generation. The current state-of-the-art text-to-audio generation methods widely adopt diffusion models (Yang et al., 2023b; Kreuk et al., 2023; Liu et al., 2023b; c; Huang et al., 2023a; Ghosal et al., 2023; Evans et al., 2024a; Vyas et al., 2023; Kreuk et al., 2023). AudioLDM (Liu et al., 2023b) makes use of a latent diffusion model conditioned on CLAP embeddings, reducing the need for the textual modality at training time. AudioLDM 2 (Liu et al., 2023c) introduces a general representation of audio unifying the tasks of music, speech, and sound effects generation. Similarly, Audiobox (Vyas et al., 2023) generates audio across different modalities such as speech and sound effects. StableAudio (Evans et al., 2024a) introduces timing embeddings to allow the generation of long audios up to 95s. Recent work also explored controllable audio generation (Shi et al., 2023; Xu et al., 2024; Melechovsky et al., 2024; Paissan et al., 2024; Zhang et al., 2023b; Liang et al., 2024; Liu et al., 2023a), visual-conditioned audio generation (Wang et al., 2024c; Mei et al., 2023b; Wang et al., 2023a), and more recently joint audio-video generation (Tang et al., 2023a; b; Xing et al., 2024; Hayakawa et al., 2024; Tian et al., 2024; Vahdati et al., 2024; Chen et al., 2024a; Kim et al., 2024a; Wang et al., 2024a; Mao et al., 2024; Chen et al., 2024c). In this work, we show that improvements to data captioning quality and size, and the adoption of scalable architecture designs lead to state-of-the-art text-to-audio generation performance.
@@
Text-Audio Datasets. The performance of text-audio models (Zhu et al., 2024; Li et al., 2023b; Deshmukh et al., 2024a; Mahfuz et al., 2023; Deshmukh et al., 2024b; Shu et al., 2023; Elizalde et al., 2024; Liu et al., 2023f; Tang et al., 2024; Gong et al., 2024; Cheng et al., 2024; Zhang et al., 2023a), including AAC, is currently hindered by the lack of high-quality large-scale paired audio text data of ambient sounds. The two main existing datasets are AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), comprising only $46k$ and $6k$ respectively of human-captioned audio clips. Another challenge is the limited availability of audio clips from sound-only platforms. LAION-Audio (Wu et al., 2023) relied on numerous sources of audio platforms such as BBC Sound Effects (BBC Sound Effects, 2024),  (Font et al., 2013) FreeSounds, and SoundBible (SoundBible, 2024) to form a dataset consisting of 630k audio samples with highly noisy raw descriptions. WavCaps (Mei et al., 2023a) proposes a filtering procedure based on ChatGPT (Achiam et al., 2023) to collect $400k$ audio clips and weakly caption them based on the noisy descriptions alone. Several subsequent work (Majumder et al., 2024; Sun et al., 2024) adopted similar strategies of using large language models to augment captions. While weak-captioning does improve downstream metrics, it is suboptimal because it fails to incorporate the audio signal itself. A recent work (Huang et al., 2023b) explored a knowledge distillation approach that leverages data labels and a pre-trained audio captioner and retriever to improve caption quality. Chen et al. (2020) attempted to extract audio clips from videos by employing classifiers to detect ambient audio, speech, and music. In this work, we introduce an efficient dataset collection pipeline that relies on video datasets to extract ambient audio clips. We use this approach to collect $57M$ audio clips and use our state-of-the-art captioning method to add audio-aligned text descriptions, compromising the largest available text-audio-video dataset.
Figure 1: (Left) Overview of AutoCap. We employ frozen CLAP (Kim et al., 2024b) and HTSAT (Chen et al., 2022) encoders to produce the audio representation. We then compact this representation into 4x fewer tokens using a Q-Former (Li et al., 2023a) module. This representation, along with tokens derived from pertinent metadata, is processed by a pretrained BART encoder-decoder model to generate the final caption. (Right) Overview of GenAu. Following latent diffusion models, we use a frozen 1D-VAE to convert a Mel-Spectrogram into latent sequences, which are then divided into groups and processed using ‘local’ attention layers based on the FIT architecture (Chen & Li, 2023). ‘Read’ and ‘write’ layers, implemented as cross-attention, facilitate information transfer between input latents and learnable latent tokens. Finally, ‘global’ attention layers on latent tokens allow for global communication across all groups.

### v3

# Taming Data and Transformers for Scalable Audio Generation

@@ Abstract
The scalability of ambient sound generators is hindered by data scarcity, insufficient caption quality, and limited scalability in model architecture. This work addresses these challenges by advancing both data and model scaling. First, we propose an efficient and scalable dataset collection pipeline tailored for ambient audio generation, resulting in AutoReCap-XL, the largest ambient audio-text dataset with over 47 million clips. To provide high-quality textual annotations, we propose AutoCap, a high-quality automatic audio captioning model. By adopting a Q-Former module and leveraging audio metadata, AutoCap substantially enhances caption quality, reaching a CIDEr score of $83.2$, a $3.2\%$ improvement over previous captioning models. Finally, we propose GenAu, a scalable transformer-based audio generation architecture that we scale up to 1.25B parameters. We demonstrate its benefits from data scaling with synthetic captions as well as model size scaling. When compared to baseline audio generators trained at similar size and data scale, GenAu obtains significant improvements of $4.7\%$ in FAD score, $11.1\%$ in IS, and $13.5\%$ in CLAP score. Our code, model checkpoints, and dataset are publicly available.
Moayed Haji-Ali1,2,∗Willi Menapace2Aliaksandr Siarohin2
Guha Balakrishnan1Sergey Tulyakov2Vicente Ordonez1
[…]
Project Webpage: https://snap-research.github.io/GenAU
@@ 1 Introduction
Text-conditioned generative models have revolutionized the field of content creation, enabling the generation of high-quality natural images (Ramesh et al., 2022; Rombach et al., 2022; Podell et al., 2023; Haji-Ali et al., 2024a), vivid videos (Ho et al., 2022; Villegas et al., 2022; Wang et al., 2024c; Qiu et al., 2023; Menapace et al., 2024), and intricate 3D shapes (Cheng et al., 2023). The domain of audio synthesis has undergone comparable advancement (Huang et al., 2023b, a; Liu et al., 2023a; Xue et al., 2024; Guan et al., 2024; Saito et al., 2024; Niu et al., 2024; Yang et al., 2023a; Evans et al., 2024a; Liu et al., 2024b; Wang et al., 2024d; Guo et al., 2024), with three broad areas of study: speech, music and ambient sounds. The success in these domains rests on two key pillars: (i) the availability of high-quality large-scale datasets with text annotations, and (ii) the development of scalable generative models (Ho et al., 2020; Song et al., 2021). The objective of this work is to improve audio generation quality by scaling ambient audio generators across both the data and model axes.
@@
In the field of audio synthesis, ambient audio generation emerges as a critical domain. Unlike speech and music, ambient sound generation is particularly challenging due to the lack of extensive, well-annotated datasets (Kim et al., 2019; Drossos et al., 2020). Attempts to curate ambient audio from online videos predominantly failed, primarily due to the dominance of speech and music content in such videos. For instance, AudioSet (Gemmeke et al., 2017), the largest available audio dataset sourced from online videos, contains $99\%$ speech or music clips. Previous efforts to filter ambient audio from similar datasets involved using expensive classifiers on the video or audio content, making it impractical to compile a large-scale dataset due to the high filtering rate. In this work, we propose a simple, yet scalable filtering approach that leverages existing automatic video transcription to identify segments with ambient sounds. This method is not only more efficient but also more feasible, as it eliminates the need to download audio or video content. Through this approach, we built AutoReCap-XL, a dataset containing 47 million ambient audio clips sourced from existing video datasets, representing a 75-fold increase over the size of previously largest available datasets.
@@
Another challenge in compiling large-scale text-audio datasets is providing accurate textual descriptions. For visual modalities, such as images and videos (Xue et al., 2022; Miech et al., 2019), researchers often relied on the raw description and metadata to train strong visual-text models including reliable captioners (Chen et al., 2024b). For ambient sounds, however, the task is substantially more challenging as accompanying raw text tends to describe visual information or convey feelings, rather than detailing the audio content. Moreover, human-captioned audio datasets are limited, containing fewer than $51k$ text-audio pairs in total. This significantly impacts the training of current captioning models, making them more susceptible to overfitting. To address this, we introduce AutoCap, a high-quality audio captioner that leverages visual cues to enhance captioning.
@@
AutoCap refines the commonly used encoder-decoder design based on a pretrained BART (Lewis et al., 2020) by introducing a Q-Former (Li et al., 2023) module which learns an intermediate representation that better aligns the encoded audio and the original BART token representation. Second, we propose to remedy the data scarcity problem by using metadata and visual cues to aid the captioning process. Critically, we augment the encoder inputs with a set of descriptive textual metadata including audio title and a caption derived from the visual modality. This dual-input approach achieves state-of-the-art performance on AudioCaps (Kim et al., 2019), marking a 3.2% improvement in CIDEr score. Using AutoCap, we provide textual descriptions for AutoReCap-XL and demonstrate the benefits of scaling audio generative models with synthetic captions.
@@
Another axis for scaling generative models is model size (Peebles & Xie, 2023). While scaling diffusion backbones has shown significant benefits in image and video generation, ambient audio generation has shown poor scaling behavior. For instance, AudioLDM2 (Liu et al., 2023b), reported worse metrics for their largest model compared to the smaller one. Similarly, EzAudio (Hai et al., 2024), achieved only marginal improvements with scaling model size. In this work, we introduce GenAu, a scalable transformer-based architecture that achieves significant improvements over state-of-the-art models. We recognize that audio grows fast temporally, yet contains many silent and redundant segments. Therefore, an efficient architecture that can handle such properties is needed. In particular, we employ a transformer architecture in the denoising backbone where we modify the FIT transformer (Chen & Li, 2023) to generate audio in the latent space. On AutoCaps dataset, GenAu achieves significant improvements over baselines, with $11.1\%$ higher Inception Score, $4.7\%$ better FAD, and $13.5\%$ improvement in CLAP score, demonstrating superior audio-text alignment and generation quality. Moreover, GenAu shows promising scaling properties, with consistent improvements across all metrics as model size increases.
@@
In summary, this work presents significant contributions in three areas: (i) AutoCap, a novel state-of-the-art audio captioner tailored towards the annotation of data at a large scale that uses audio metadata to improve accuracy and robustness; (ii) AutoReCap-XL, a large scale ambient audio dataset, comprising 47M audio clips paired with synthetic captions, 75 times larger than available datasets (iii) GenAu, a novel audio generator based on a scalable transformer architecture specifically adapted to the audio domain, achieving significant improvements over previous state-of-the-art.
Figure 1: (Left) Overview of AutoCap. We employ a frozen HTSAT (Chen et al., 2022) encoder to produce an audio fine-grained representation of 1024 tokens. We then employ a Q-Former (Li et al., 2023) module to produce 256 tokens. These tokens, along with audio CLAP embeddings (Wu et al., 2023a) and 64 tokens of pertinent metadata, are processed by a pretrained BART to generate the final caption. (Right) Overview of GenAu. Following latent diffusion models, we use a frozen 1D-VAE to convert a Mel-Spectrogram into latent sequences, which are then divided into groups and processed using ‘local’ attention layers based on the FIT architecture (Chen & Li, 2023). ‘Read’ and ‘write’ layers, implemented as cross-attention, facilitate information transfer between input latents and learnable latent tokens. Finally, ‘global’ attention layers on latent tokens facilitate global communication across groups.
@@ 2 Related Work
Automatic Audio Captioning (AAC). The goal of AAC is to produce language descriptions for given audio content. Recent AAC methods (Deshmukh et al., 2024a; Gontier et al., 2021; Wu et al., 2024; Salewski et al., 2023; Sridhar et al., 2024; Kadlčík et al., 2023; Cousin et al., 2023; Labbé et al., 2023; Xu et al., 2024b; Zhang et al., 2024; Ghosh et al., 2024a; Deshmukh et al., 2024a) employ encoder-decoder transformer architectures, where an encoder receiving the audio signal produces a representation that is used by the decoder to produce the output caption. WavCaps (Mei et al., 2024a) employs HTSAT (Chen et al., 2022) audio encoder and uses a pretrained BART (Lewis et al., 2020) as the decoder. Similarly, EnCLAP (Kim et al., 2024b) uses a pretrained BART and improves on the audio representation. Recognizing the limited data, CoNeTTE (Labb et al., 2024) proposes to train a lightweight vanilla transformer decoder (Vaswani et al., 2017) instead. Other work explored augmentation to counter data scarcity (Kim et al., 2022; Labb et al., 2024; Ye et al., 2022). Recent work (Liu et al., 2023c; Sun et al., 2024; Yuan et al., 2024) proposed to leverage visual information to address sound ambiguities, reporting improvements. More recent methods use audio large language model for zero-shot captioning (Kong et al., 2024; Deshmukh et al., 2023; Ghosh et al., 2024b). Our method uses audio metadata and visual information as additional signals and leverages a lightweight Q-Former (Li et al., 2023) model to improve accuracy.
@@
Text-conditioned audio generation. The current state-of-the-art text-to-audio generation methods widely adopt diffusion models (Yang et al., 2023b; Kreuk et al., 2023; Liu et al., 2023a, b; Huang et al., 2023a; Ghosal et al., 2023; Evans et al., 2024a; Vyas et al., 2023; Kreuk et al., 2023; Hai et al., 2024; Evans et al., 2024b). AudioLDM 1 & 2 (Liu et al., 2023a, b) make use of a latent diffusion model and employ a UNet as the diffusion backbone. Recently, StableAudio Open (Evans et al., 2024b) introduced a 1.32B model that uses a DiT (Peebles & Xie, 2023) to generate variable-length audio clips at 48kHz. Recent work also explored controllable audio generation (Shi et al., 2023; Xu et al., 2024a; Melechovsky et al., 2023; Paissan et al., 2023; Zhang et al., 2023b; Liang et al., 2024; Liu et al., 2024a), visual-conditioned audio generation (Wang et al., 2024d; Mei et al., 2024b; Wang et al., 2024a), and more recently joint audio-video generation (Tang et al., 2024, 2023b; Xing et al., 2024; Hayakawa et al., 2024; Tian et al., 2024; Vahdati et al., 2024; Chen et al., 2024a; Kim et al., 2024a; Wang et al., 2024b; Mao et al., 2024; Haji-Ali et al., 2024b). In this work, we propose a transformer architecture design that shows strong scalability properties.
@@
Text-Audio Datasets. The performance of text-audio models (Zhu et al., 2024; Li et al., 2024; Deshmukh et al., 2023; Mahfuz et al., 2023; Deshmukh et al., 2024b; Shu et al., 2023; Elizalde et al., 2024; Liu et al., 2024c; Tang et al., 2023a; Gong et al., 2023; Cheng et al., 2024; Zhang et al., 2023a) is currently hindered by the lack of high-quality large-scale paired audio text data of ambient sounds. Existing human-captioned (AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020)), have in total only $51k$ audio-text pairs. Another challenge is the limited availability of audio clips from sound-only platforms. LAION-Audio (Wu et al., 2023a) relied on numerous sources of audio platforms such as BBC Sound Effects (BBC Sound Effects, 2024),  (Font et al., 2013) FreeSounds, and SoundBible (SoundBible, 2024) to form a dataset consisting of 630k audio samples with highly noisy raw descriptions. Chen et al. (2020) attempted to extract audio clips from videos by employing classifiers to detect ambient audio, speech, and music. To annotate these clips, WavCaps (Mei et al., 2024a) proposes a filtering procedure based on ChatGPT (Achiam et al., 2023) to collect $400k$ audio clips and weakly caption them based on the noisy descriptions alone. Several subsequent work (Majumder et al., 2024; Sun et al., 2024) adopted similar strategies of using large language models to augment captions. While weak-captioning improves downstream metrics, it is suboptimal because it fails to incorporate the audio signal itself. In this work, we introduce an efficient dataset collection pipeline that relies on video datasets to extract ambient audio clips and automatic captioners to provide textual descriptions. We collect $47M$ audio clips, marking the largest available text-audio dataset.